OpenStudy (anonymous):

Use this to find the equation of the tangent line to the parabola y=3x 2 −5x+5 at the point (3,17) . The equation of this tangent line can be written in the form y=mx+b 1. What is (m) Slope? 2. What is (b) y-intercept ?

OpenStudy (anonymous):

ok First thing to do is to find the derivative of \[3x^2 -5x + 5\] Do you know how to do that?

OpenStudy (anonymous):

6x-5

OpenStudy (anonymous):

Excellent! Now plug in the x value of the point (3, 17) into \[6x - 5\] and that is how you find the slope. So what do you get when you do that?

OpenStudy (anonymous):

13 !

OpenStudy (anonymous):

and then y intercept is -22 ! yayyyy

OpenStudy (anonymous):

Good. Now you have a slope of m = 13 and a point of (3,17). So using the point-slope formula which is \[y - y{1} = m(x - x{1})\] plugging in those values you get \[y - 17 = 13(x - 3)\] solve for y and that is the equation of the tangent line. so \[y = 13x - 22\]

OpenStudy (anonymous):

And you are correct! Nicely done!